OpenStudy (anonymous):

which of the following represents a function? x=y^2 x^2 + y^2 =49 x=[[y]] y=absolute x

OpenStudy (anonymous):

help please n_n urgent

OpenStudy (acxbox22):

a function represents y as a function of x which answer choice demonstrates that?

OpenStudy (anonymous):

x=[[y]]

OpenStudy (anonymous):

the rest have more than one intersection

OpenStudy (acxbox22):

hint- a function starts with a y= for example: y=x, y=x^2,y=sqrt x

OpenStudy (anonymous):

mmmhhh is my reasoning wrong

OpenStudy (anonymous):

ohhhh the absolute one !! @acxbox22

OpenStudy (anonymous):

@phi would this be correct ?

OpenStudy (acxbox22):



OpenStudy (anonymous):

really wow i thought it would be the line thank you @acxbox22

OpenStudy (anonymous):

It depends on what you call the independent variable. \(x=y^2\) is a function of \(x\) in terms of \(y\), as well as \(x=[[y]]\).

OpenStudy (acxbox22):

generally the independent variable is x